# Twilight: Adaptive Attention Sparsity with Hierarchical Top-$p$ Pruning

### v3

@@ 1 Introduction
Large language models (LLMs) with long-context capabilities have revolutionized a wide array of natural language processing applications, such as retrieval-based tasks, document summarization bai2024longbench , and code generation jain2024livecodebenchholisticcontaminationfree . The increasing availability of models supporting context windows up to 1M to 10M tokens yang2025qwen251mtechnicalreport ; geminiteam2024gemini15unlockingmultimodal highlights the growing potential of these advancements. For instance, video language models (VLMs) wang2024qwen2vlenhancingvisionlanguagemodels often require tens of thousands of tokens for video processing. Similarly, large reasoning models deepseekai2025deepseekr1incentivizingreasoningcapability ; kimiteam2025kimik15scalingreinforcement , which are rapidly growing in popularity, frequently demand substantial token lengths to enable chain-of-thought (CoT) reasoning. Consequently, the importance of long-context LLMs is increasing rapidly to meet the needs of these sophisticated applications.
@@
Despite the substantial potential of long-context LLMs, they come with excessive computational and memory costs zhang2023h2o ; tang2024quest ; xiao2024duo , primarily from the attention mechanism. Particularly, in the decoding stage of LLMs, the key-value (KV) cache size grows rapidly as the token sequence becomes longer. These data need to be repeatedly loaded from the memory, leading to significant latency overheads. Furthermore, the substantial size of the KV cache significantly increases the GPU memory consumption, compounding the challenges of continuously scaling long-context LLMs.
Previous research has extensively investigated the use of attention sparsity (a.k.a., KV cache sparsity) to accelerate long-context inference, both during the prefilling and decoding stages. The core idea is to compute an approximate attention on a subset of tokens, often referred to as “critical tokens” or “heavy hitters” zhang2023h2o . The number of selected tokens, denoted as $B$, is commonly referred to as the KV cache budget. A top-$k$ operation is required to identify the indices of the critical tokens that correspond to the $B$ highest estimated scores. However, a key tradeoff exists for the selection of the budget. A smaller $B$ value significantly reduces the memory accesses and computations, while a larger $B$ value retains more contextual information and thereby minimizes the accuracy loss.
@@
Identifying the optimal value of $B$ to balance both accuracy and efficiency is inherently challenging due to two major reasons: (a) The best budget choices vary dynamically at runtime. Previous works wu2024retrievalheadmechanisticallyexplains ; xiao2024duo have demonstrated that some heads, referred to as “retrieval heads”, are trained to extract important information from long contexts, while others focus only on local information. From Figure 1 we see that the distribution of attention weights may vary across different attention heads. Some attention distributions concentrate on a small subset of tokens, which we refer to as focused attention. Other attention distributions may be flatter, where many tokens have similar attention weights; we call this diffuse attention. For focused attention, using a fixed token budget for top-$k$ attention often leads to over-selection, as only a few tokens are needed to accumulate sufficient attention weights. In contrast, for diffuse attention, a fixed budget can result in under-selection, as a larger number of tokens are necessary to ensure accurate attention modeling. (b) Existing algorithms need different degrees of over-selection to compensate the estimation inaccuracy. As shown in Figure 2, the optimized budgets highly depend on the specific algorithms, necessitating offline calibration to determine the appropriate budget for each algorithm individually. Certain methods, like Quest tang2024quest or DS yang2024post , have to over-select some tokens to compensate for the inevitable inaccuracy in estimating the importance of tokens compared to the oracle.
@@
Our evaluations for Twilight are conducted in two aspects: accuracy and efficiency. First, we demonstrate that Twilight optimizes the base algorithms with nearly no accuracy loss on both medium-context benchmarks (GSM8K cobbe2021training , COQA reddy2019coqa , PG-19 rae2019compressive ) and long-context benchmarks (Longbench bai2024longbench , RULER hsieh2024ruler ). Next, we show that Twilight achieves up to $15.8\times$ performance improvement over the full attention operation. Compared to prior sparse attention methods, Twilight enables a $1.4\times$ speedup for the self-attention operator itself, and a $1.35\times$ speedup for the end-to-end decoding. Our contributions are summarized as follows:
@@ 2 Related Work
Top-$k$ Sparse Attention. H2O zhang2023h2o , StreamingLLM xiao2023streamingllm , and SnapKV li2024snapkv evict non-critical tokens in static, query-agnostic manners, which are often referred to as KV cache compression. In contrast, SparQ ribar2023sparq , Quest tang2024quest , Double Sparsity (DS) yang2024post , and HShare wu2025hshare retain all tokens in the GPU memory but select critical tokens to save data accesses. Recent works like RetrievalAttention liu2024retrievalattention and PQCache zhang2025pqcache adopt advanced algorithms to better estimate the token criticality. NSA yuan2025native and MoBA lu2025moba explore opportunities in trainable sparse attention. However, these methods are all based on top-$k$ which requires proper budget selection and configuration beforehand, and thus suffer from the over/under-selection issues.
Dynamic Budget. More recent studies have extensively demonstrated that the optimal budgets vary significantly across different layers cai2024pyramidkv ; yang2024pyramidinfer , attention heads feng2025adakvoptimizingkvcache ; xiao2024duo ; tang2024razorattention , and prompts (tasks) zhou2024dynamickv . Please see Appendix A for details. These works tend to focus on only one aspect of the dynamism. In this paper, we point out that it is the different distributions of attention weights that are the root cause of such dynamism.
Non-top-$k$ Sparse Attention. Some recently emerged designs also go beyond top-$k$ methods. MagicPIG chen2024magicpig uses locality-sensitive hash (LSH) sampling instead of dropping tokens to estimate attention weights, but requires complicated algorithm-system co-design. SampleAttention zhu2024sampleattention also features adaptive sparsity but focuses on the prefill stage. A concurrent work with ours, Tactic zhu2025tactic , also dives into top-$p$ sparsity but it uses function fitting to estimate the weight distributions. Although it potentially has lower estimation cost, it usually overestimates the budget.
Other KV Cache Optimizations. Several alternative approaches focus on optimizing the KV cache beyond sparsification, including quantization hooper2024kvquant10millioncontext ; liu2024kivi ; kang2024gearefficientkvcache ; nawrot2024dynamicmemorycompressionretrofitting , linear attention wang2020linformerselfattentionlinearcomplexity ; katharopoulos2020transformersrnnsfastautoregressive , and memory-efficient attention mechanisms such as FlashAttention dao2023flashattention2 and SageAttention zhang2025sageattention ; zhang2024sageattention2 . Our approach is orthogonal to these methods, and can be combined with them for enhanced performance.
@@ 3.1 Problem Formulation
Let the accurate attention output be $\mathbf{o}=\text{softmax}(\frac{\mathbf{q}\cdot\mathbf{K}^{T}}{\sqrt{d}})\mathbf{V}\in\mathbb{R}^{1\times d}$. To minimize the error $\|\mathbf{o}-\hat{\mathbf{o}}\|$, we need to carefully select the subset of tokens used in $\mathcal{I}$. However, directly optimizing this objective function without loading the full KV cache is challenging. According to the sub-multiplicative property of the Frobenius norm, we can bound the error as in Equation 2. Earlier research has shown that the distribution of $\mathbf{V}$ is relatively smooth zhao2024atom , which implies $\|\mathbf{V}\|_{F}$ can be viewed as a constant.
@@ 3.2 Rethinking the Problem of Top-$k$
The Achilles’ heel of top-$k$ sparse attention, as described earlier, is the dilemma in determining a universally applicable budget $B$ to all scenarios. We find that this predicament is quite similar to a previous problem encountered in the sampling phase of LLMs, during which the model samples the final output token from the predicted probability distribution. Nucleus sampling holtzman2019curious , a.k.a., top-$p$ sampling, was proposed to address the problem that top-$k$ sampling cannot adapt to different next-word distributions.
@@ 4 Twilight
With the efficient and adaptive top-$p$ sparse attention, our primary goal is to use it to endow existing algorithms with adaptive budget selection capabilities, rather than simply inventing yet another sparse attention design. We are mainly motivated by two reasons. On one hand, despite the challenge of budget selection, existing sparse attention algorithms have achieved significant success in LLM serving systems kwon2023vllm ; zheng2023sglang , thanks to their effective token selection strategies. These strategies can be readily reused and enhanced with our adaptive sparsity. On the other hand, we anticipate that future sparse attention methods may still employ top-$k$ selection. By developing a general solution, we aim to automatically equip these future methods with adaptive attention sparsity, while avoiding extensive redesign. Consequently, we position our system, Twilight, as an optimizer for existing algorithms.
@@
Nevertheless, applying top-$p$ to various existing sparse attention algorithms faces three key challenges on both the algorithm and system perspectives. (C1) Not all algorithms are suitable for top-$p$. Top-$p$ imposes strict constraints on the layout of attention weights. For example, simply replacing top-$k$ with top-$p$ in Quest tang2024quest would not work, as Quest performs max pooling on weights with a per-page layout (16 tokens per page). Additionally, some other methods yang2024tidaldecode ; liu2024retrievalattention do not use attention weights to select critical tokens at all. (C2) It is harder to estimate weights for top-$p$ than top-$k$. In order to find critical tokens without loading full K data, low-precision representation of the K cache is usually used yang2024post ; zhang2025pqcache . However, the precision requirement of top-$p$ is higher than that of top-$k$, because the former requires a certain degree of numerical accuracy while the latter only demands ordinality. Table 1 compares top-$k$, top-$p$, and full attention. The precision requirement of top-$p$ attention lies in between the other two, necessitating reconsideration of appropriate precision choices for the K cache. (C3) System-level optimizations are needed. Since our work is the first to introduce top-$p$ to attention weights, the relevant algorithms need to be efficiently implemented on the GPU, including efforts on both parallel algorithm designs and kernel optimizations.
@@ 4.2 Efficient Kernel Implementations
Efficient SpGEMV with 4-bit Quantization of Key Cache. The beginning part of the Pruner is similar to other sparse attention algorithms, which is to estimate the importance of tokens. As we formulated in Section 3.1, this can be done by estimating the similarity between $\mathbf{q}$ and $\mathbf{K}$, i.e., $\mathbf{q}\cdot\mathbf{K}$. Since loading $\mathbf{K}$ is known to be memory bound, we reduce the memory access cost by quantizing $\mathbf{K}$ into lower precision. But what precision shall we choose? Table 1 shows the precision requirement of top-$p$ lies in between top-$k$ and full attention. Some existing top-$k$ designs yang2024post ; zhang2025pqcache have pushed the compression of the K cache to extremely low precisions of 1 to 2 bits. For full attention, SageAttention zhang2025sageattention has demonstrated 8-bit precision with smoothing $\mathbf{K}$ and per-block quantization. In this work, we empirically find that 4-bit precision strikes a balance between accuracy and efficiency for top-$p$, as illustrated in Figure 6. Here the sum of attention weights with 2-bit quantization drops significantly, while 4-bit and 8-bit methods both maintain enough stability.
Hence we implement an efficient sparse GEMV (SpGEMV) kernel based on FlashInfer ye2025flashinfer , a high-performance kernel library for LLM serving. Here “sparse” means the quantized K cache data are stored/loaded in a paged manner kwon2023vllm to align with the original KV cache layout. We maintain this extra INT4 asymmetrically quantized K cache on the GPU as shown at the right of Figure 5. The INT4 $\mathbf{K}$ vectors are unpacked and dequantized in the shared memory, reducing data accesses from the global memory to at most $1/4$, resulting in considerable end-to-end speedup.
@@
Efficient Top-$p$ via Binary Search. A brute-force way to do top-$p$ sampling is to sort the elements by descending order and accumulate them until the sum meets the threshold. This is quite inefficient in parallel hardware like modern GPUs. As our top-$p$ method is motivated by the top-$p$ sampling, we also implement this kernel by modifying the top-$p$ sampling kernel from FlashInfer ye2025flashinfer . Specifically, our kernel adopts a parallel-friendly binary search method as in Algorithm 1. Note that element-wise operations like max, where, and sum can be fused into a single loop, which is tensorized on the GPU. Thus we do not need to materialize the intermediate variables like $W_{0}$.
Load Balancing with Awareness of Head Dynamism. The top-$p$ Pruner enables head-wise dynamic budgets, but also raises load imbalance issues in the attention kernel. Traditional implementations allocate uniform computation resources to all heads. FlashInfer ye2025flashinfer deeply investigates this load imbalance problem, but only for requests with dynamic lengths. Twilight further reuses the load balancing algorithm in FlashInfer to address head-wise dynamism, by flattening the head dimension.
@@ 4.3 Overhead Analysis and Discussion
Memory Overheads. Twilight introduces an extra INT4 quantized K cache, which brings a ${1}/{2}\times{1}/{4}={1}/{8}$ extra KV cache memory overhead. However, this additional cost does not appear in all cases. First, some base algorithms, like DS yang2024post , already maintain an INT4 K cache. Second, some recent efforts have explored INT4 full attention zhang2024sageattention2 . This allows us to directly reuse the estimated attention weights calculated by the INT4 K cache in the attention computation, without maintaining the original FP16 K cache. Moreover, offloading and selective quantization (e.g., keeping the extra INT4 K cache only for hot tokens) can be leveraged if the GPU memory becomes a bottleneck, which we leave as future work.
Integration with LLM Serving Systems. Our system design naturally aligns with PagedAttention kwon2023vllm , so Twilight can be seamlessly integrated into popular serving systems like vLLM kwon2023vllm and SGLang zheng2023sglang . Other common techniques, such as prefix sharing and multi-phase attention lin2024parrot ; zheng2023sglang ; zhu2024relayattention ; ye2024chunkattention ; ye2024cascade , are also compatible with Twilight since we use page-level or token-level sparse operations, and can achieve a flexible computation flow.

### v5

@@ 1 Introduction
Large language models (LLMs) with long-context capabilities have revolutionized a wide array of natural language processing applications, such as retrieval-based tasks, document summarization [1], and code generation [2]. The increasing availability of models supporting context windows up to 1M to 10M tokens [3, 4] highlights the growing potential of these advancements. For instance, video language models (VLMs) [5] often require tens of thousands of tokens for video processing. Similarly, large reasoning models [6, 7], which are rapidly growing in popularity, frequently demand substantial token lengths to enable chain-of-thought (CoT) reasoning. Consequently, the importance of long-context LLMs is increasing rapidly to meet the needs of these sophisticated applications.
@@
Despite the substantial potential of long-context LLMs, they come with excessive computational and memory costs [8, 9, 10], primarily from the attention mechanism. Particularly, in the decoding stage of LLMs, the key-value (KV) cache size grows rapidly as the token sequence becomes longer. These data need to be repeatedly loaded from the memory, leading to significant latency overheads. Furthermore, the substantial size of the KV cache significantly increases the GPU memory consumption, compounding the challenges of continuously scaling long-context LLMs.
Previous research has extensively investigated the use of attention sparsity (a.k.a., KV cache sparsity) to accelerate long-context inference, both during the prefilling and decoding stages. The core idea is to compute an approximate attention on a subset of tokens, often referred to as “critical tokens” or “heavy hitters” [8]. The number of selected tokens, denoted as $B$, is commonly referred to as the KV cache budget. A top-$k$ operation is required to identify the indices of the critical tokens that correspond to the $B$ highest estimated scores. However, a key tradeoff exists for the selection of the budget. A smaller $B$ value significantly reduces the memory accesses and computations, while a larger $B$ value retains more contextual information and thereby minimizes the accuracy loss.
@@
Identifying the optimal value of $B$ to balance both accuracy and efficiency is inherently challenging due to two major reasons: (a) The best budget choices vary dynamically at runtime. Previous works [11, 10] have demonstrated that some heads, referred to as “retrieval heads”, are trained to extract important information from long contexts, while others focus only on local information. From Figure 1 we see that the distribution of attention weights may vary across different attention heads. Some attention distributions concentrate on a small subset of tokens, which we refer to as focused attention. Other attention distributions may be flatter, where many tokens have similar attention weights; we call this diffuse attention. For focused attention, using a fixed token budget for top-$k$ attention often leads to over-selection, as only a few tokens are needed to accumulate sufficient attention weights. In contrast, for diffuse attention, a fixed budget can result in under-selection, as a larger number of tokens are necessary to ensure accurate attention modeling. (b) Existing algorithms need different degrees of over-selection to compensate the estimation inaccuracy. As shown in Figure 2, the optimized budgets highly depend on the specific algorithms, necessitating offline calibration to determine the appropriate budget for each algorithm individually. Certain methods, like Quest [9] or DS [12], have to over-select some tokens to compensate for the inevitable inaccuracy in estimating the importance of tokens compared to the oracle.
@@
Our evaluations for Twilight are conducted in two aspects: accuracy and efficiency. First, we demonstrate that Twilight optimizes the base algorithms with nearly no accuracy loss on both medium-context benchmarks (GSM8K [13], COQA [14], PG-19 [15]) and long-context benchmarks (Longbench [1], RULER [16]). Next, we show that Twilight achieves up to $15.8\times$ performance improvement over the full attention operation. Compared to prior sparse attention methods, Twilight enables a $1.4\times$ speedup for the self-attention operator itself, and a $1.35\times$ speedup for the end-to-end decoding. Our contributions are summarized as follows:
@@ 2 Related Work
Top-$k$ Sparse Attention. H2O [8], StreamingLLM [17], and SnapKV [18] evict non-critical tokens in static, query-agnostic manners, which are often referred to as KV cache compression. In contrast, SparQ [19], Quest [9], Double Sparsity (DS) [12], and HShare [20] retain all tokens in the GPU memory but select critical tokens to save data accesses. Recent works like RetrievalAttention [21] and PQCache [22] adopt advanced algorithms to better estimate the token criticality. NSA [23] and MoBA [24] explore opportunities in trainable sparse attention. However, these methods are all based on top-$k$ which requires proper budget selection and configuration beforehand, and thus suffer from the over/under-selection issues.
Dynamic Budget. More recent studies have extensively demonstrated that the optimal budgets vary significantly across different layers [25, 26], attention heads [27, 10, 28], and prompts (tasks) [29]. Please see Appendix A for details. These works tend to focus on only one aspect of the dynamism. In this paper, we point out that it is the different distributions of attention weights that are the root cause of such dynamism.
Non-top-$k$ Sparse Attention. Some recently emerged designs also go beyond top-$k$ methods. MagicPIG [30] uses locality-sensitive hash (LSH) sampling instead of dropping tokens to estimate attention weights, but requires complicated algorithm-system co-design. SampleAttention [31] also features adaptive sparsity but focuses on the prefill stage. A concurrent work with ours, Tactic [32], also dives into top-$p$ sparsity but it uses function fitting to estimate the weight distributions. Although it potentially has lower estimation cost, it usually overestimates the budget.
Other KV Cache Optimizations. Several alternative approaches focus on optimizing the KV cache beyond sparsification, including quantization [33, 34, 35, 36], linear attention [37, 38], and memory-efficient attention mechanisms such as FlashAttention [39] and SageAttention [40, 41]. Our approach is orthogonal to these methods, and can be combined with them for enhanced performance.
@@ 3.1 Problem Formulation
Let the accurate attention output be $\mathbf{o}=\text{softmax}(\frac{\mathbf{q}\cdot\mathbf{K}^{T}}{\sqrt{d}})\mathbf{V}\in\mathbb{R}^{1\times d}$. To minimize the error $\|\mathbf{o}-\hat{\mathbf{o}}\|$, we need to carefully select the subset of tokens used in $\mathcal{I}$. However, directly optimizing this objective function without loading the full KV cache is challenging. According to the sub-multiplicative property of the Frobenius norm, we can bound the error as in Equation 2. Earlier research has shown that the distribution of $\mathbf{V}$ is relatively smooth [42], which implies $\|\mathbf{V}\|_{F}$ can be viewed as a constant.
@@ 3.2 Rethinking the Problem of Top-$k$
The Achilles’ heel of top-$k$ sparse attention, as described earlier, is the dilemma in determining a universally applicable budget $B$ to all scenarios. We find that this predicament is quite similar to a previous problem encountered in the sampling phase of LLMs, during which the model samples the final output token from the predicted probability distribution. Nucleus sampling [43], a.k.a., top-$p$ sampling, was proposed to address the problem that top-$k$ sampling cannot adapt to different next-word distributions.
@@ 4 Twilight
With the efficient and adaptive top-$p$ sparse attention, our primary goal is to use it to endow existing algorithms with adaptive budget selection capabilities, rather than simply inventing yet another sparse attention design. We are mainly motivated by two reasons. On one hand, despite the challenge of budget selection, existing sparse attention algorithms have achieved significant success in LLM serving systems [44, 45], thanks to their effective token selection strategies. These strategies can be readily reused and enhanced with our adaptive sparsity. On the other hand, we anticipate that future sparse attention methods may still employ top-$k$ selection. By developing a general solution, we aim to automatically equip these future methods with adaptive attention sparsity, while avoiding extensive redesign. Consequently, we position our system, Twilight, as an optimizer for existing algorithms.
@@
Nevertheless, applying top-$p$ to various existing sparse attention algorithms faces three key challenges on both the algorithm and system perspectives. (C1) Not all algorithms are suitable for top-$p$. Top-$p$ imposes strict constraints on the layout of attention weights. For example, simply replacing top-$k$ with top-$p$ in Quest [9] would not work, as Quest performs max pooling on weights with a per-page layout (16 tokens per page). Additionally, some other methods [46, 21] do not use attention weights to select critical tokens at all. (C2) It is harder to estimate weights for top-$p$ than top-$k$. In order to find critical tokens without loading full K data, low-precision representation of the K cache is usually used [12, 22]. However, the precision requirement of top-$p$ is higher than that of top-$k$, because the former requires a certain degree of numerical accuracy while the latter only demands ordinality. Table 1 compares top-$k$, top-$p$, and full attention. The precision requirement of top-$p$ attention lies in between the other two, necessitating reconsideration of appropriate precision choices for the K cache. (C3) System-level optimizations are needed. Since our work is the first to introduce top-$p$ to attention weights, the relevant algorithms need to be efficiently implemented on the GPU, including efforts on both parallel algorithm designs and kernel optimizations.
@@ 4.2 Efficient Kernel Implementations
Efficient SpGEMV with 4-bit Quantization of Key Cache. The beginning part of the Pruner is similar to other sparse attention algorithms, which is to estimate the importance of tokens. As we formulated in Section 3.1, this can be done by estimating the similarity between $\mathbf{q}$ and $\mathbf{K}$, i.e., $\mathbf{q}\cdot\mathbf{K}$. Since loading $\mathbf{K}$ is known to be memory bound, we reduce the memory access cost by quantizing $\mathbf{K}$ into lower precision. But what precision shall we choose? Table 1 shows the precision requirement of top-$p$ lies in between top-$k$ and full attention. Some existing top-$k$ designs [12, 22] have pushed the compression of the K cache to extremely low precisions of 1 to 2 bits. For full attention, SageAttention [40] has demonstrated 8-bit precision with smoothing $\mathbf{K}$ and per-block quantization. In this work, we empirically find that 4-bit precision strikes a balance between accuracy and efficiency for top-$p$, as illustrated in Figure 6. Here the sum of attention weights with 2-bit quantization drops significantly, while 4-bit and 8-bit methods both maintain enough stability.
Hence we implement an efficient sparse GEMV (SpGEMV) kernel based on FlashInfer [47], a high-performance kernel library for LLM serving. Here “sparse” means the quantized K cache data are stored/loaded in a paged manner [44] to align with the original KV cache layout. We maintain this extra INT4 asymmetrically quantized K cache on the GPU as shown at the right of Figure 5. The INT4 $\mathbf{K}$ vectors are unpacked and dequantized in the shared memory, reducing data accesses from the global memory to at most $1/4$, resulting in considerable end-to-end speedup.
@@
Efficient Top-$p$ via Binary Search. A brute-force way to do top-$p$ sampling is to sort the elements by descending order and accumulate them until the sum meets the threshold. This is quite inefficient in parallel hardware like modern GPUs. As our top-$p$ method is motivated by the top-$p$ sampling, we also implement this kernel by modifying the top-$p$ sampling kernel from FlashInfer [47]. Specifically, our kernel adopts a parallel-friendly binary search method as in Algorithm 1. Note that element-wise operations like max, where, and sum can be fused into a single loop, which is tensorized on the GPU. Thus we do not need to materialize the intermediate variables like $W_{0}$.
Load Balancing with Awareness of Head Dynamism. The top-$p$ Pruner enables head-wise dynamic budgets, but also raises load imbalance issues in the attention kernel. Traditional implementations allocate uniform computation resources to all heads. FlashInfer [47] deeply investigates this load imbalance problem, but only for requests with dynamic lengths. Twilight further reuses the load balancing algorithm in FlashInfer to address head-wise dynamism, by flattening the head dimension.
@@ 4.3 Overhead Analysis and Discussion
Memory Overheads. Twilight introduces an extra INT4 quantized K cache, which brings a ${1}/{2}\times{1}/{4}={1}/{8}$ extra KV cache memory overhead. However, this additional cost does not appear in all cases. First, some base algorithms, like DS [12], already maintain an INT4 K cache. Second, some recent efforts have explored INT4 full attention [41]. This allows us to directly reuse the estimated attention weights calculated by the INT4 K cache in the attention computation, without maintaining the original FP16 K cache. Moreover, offloading and selective quantization (e.g., keeping the extra INT4 K cache only for hot tokens) can be leveraged if the GPU memory becomes a bottleneck, which we leave as future work.
Integration with LLM Serving Systems. Our system design naturally aligns with PagedAttention [44], so Twilight can be seamlessly integrated into popular serving systems like vLLM [44] and SGLang [45]. Other common techniques, such as prefix sharing and multi-phase attention [48, 45, 49, 50, 51], are also compatible with Twilight since we use page-level or token-level sparse operations, and can achieve a flexible computation flow.